# Can We Count on LLMs? The Fixed-Effect Fallacy and Claims of GPT-4 Capabilities

### v1

@@ 1 Introduction
We conclude that efforts to quantify LLM capabilities easily succumb to the language-as-fixed-effect fallacy (Clark, 1973; Coleman, 1964; Yarkoni, 2019), where experimental observations on language-tasks are improperly generalized beyond what the data supports. A consequence appears to be that intuitions that have been formed based on interactions with humans form a very unreliable guide as to which modifications should “make no difference” to LLM performance. For example, the abstractions that we take for granted for humans (e.g., of separating the task of counting from the thing-to-be-counted) do not appear to be replicated by LLMs.
@@ 4.1 Experimental setup
Unless otherwise specified all of the conditions were evaluated on $500$ independent runs. Thus, for example, if a table entry reports a success rate of 89.0% on a task, and sampling were the only source of randomness, then a reasonable estimate of the 95% confidence interval would be $1.96\cdot\sqrt{0.89\times 0.11/500}\approx 2.74\%.$ However, an important finding, below, is that there are significant other sources of randomness, and the conventional way of estimating margins-of-error cannot be applied. All of the trials are performed using the OpenAI GPT-4 API with a temperature setting of $0.7$ The results of all queries are available in the GitHub repository.
@@ 5 Related Work
Yarkoni (Yarkoni, 2019) argues that the problem of improper generalization goes far beyond the language issue. He suggests that confusing fixed effects for random ones is the source of many of the replication failures in the social sciences.
@@ 7 Conclusion
We find that even when modifications are trivial and make no difference to human performance on a task we cannot assume that the same is true of LLM performance. In the absence of evidence to the contrary, measurements of LLM task-accuracy cannot be assumed to generalize beyond the precise conditions studied.
@@ Appendix A Appendix
Here we revisit accuracy measurements for the counting task studied in Section 4.2 but using the GPT-3.5, Mistral Instruct 7B Q4 and Llama 3 8B Q4 models. GPT-3.5 was accessed via the openai API. The Mistral and Llama models were run locally using versions with quantized coefficients. Each cell in each table represents $500$ trials. All trials were done using a temperature setting of $0.7$

### v2

@@ 1 Introduction
We conclude that efforts to quantify LLM capabilities easily succumb to the language-as-fixed-effect fallacy (Clark, 1973; Coleman, 1964; Yarkoni, 2022), where experimental observations on language-tasks are improperly generalized beyond what the data supports. A consequence appears to be that intuitions that have been formed based on interactions with humans form a very unreliable guide as to which modifications should “make no difference” to LLM performance. For example, the abstractions that we take for granted for humans (e.g., of separating the task of counting from the thing-to-be-counted) do not appear to be replicated by LLMs.
@@ 4.1 Experimental setup
Unless otherwise specified all of the conditions were evaluated on $500$ independent runs. Thus, for example, if a table entry reports a success rate of 89.0% on a task, and sampling were the only source of randomness, then a reasonable estimate of the 95% confidence interval would be $1.96\cdot\sqrt{0.89\times 0.11/500}\approx 2.74\%.$ However, an important finding, below, is that there are significant other sources of randomness, and the conventional way of estimating margins-of-error cannot be applied. Lists were generated independently for each trial at query-time; thus, we did not re-use lists across conditions. All of the trials are performed using the OpenAI GPT-4 API with a temperature setting of $0.7$ The results of all queries are available in the GitHub repository https://github.com/demarinaGit/canWeCountOnLLMs. All trials were performed using a temperature setting of 0.9.
@@ 5 Related Work
Yarkoni (Yarkoni, 2022) argues that the problem of improper generalization goes far beyond the language issue. He suggests that confusing fixed effects for random ones is the source of many of the replication failures in the social sciences.
@@ 7 Conclusion
We find that, even when modifications are trivial and make no difference to human performance on a task, we cannot assume that the same is true of LLM performance. In the absence of evidence to the contrary, measurements of LLM task-accuracy cannot be assumed to generalize beyond the precise conditions studied.
@@ Appendix A Appendix
Here we revisit accuracy measurements for the counting task studied in Section 4.2 but using the GPT-3.5, Mistral Instruct 7B Q4 and Llama 3 8B Q4 models. GPT-3.5 was accessed via the openai API. The Mistral and Llama models were run locally using versions with quantized coefficients. Each cell in each table represents $500$ trials.